# Adaptively Robust and Sparse

### v1

@@ Introduction
When applying the $P_{2}(w_{j};\lambda_{2})=\lambda_{2}|w_{j}|$ penalty in (9), the superscript $S$ denotes the soft-thresholding operator as:
@@
When applying the $P_{2}(w_{j};\lambda_{2})=P^{SCAD}_{\lambda_{2}}(w_{j})$ penalty in (9), the superscript $S$ denotes SCAD-thresholding operator as:
@@
The parameter $\lambda_{2}$ has a crucial role in controlling the sparsity in the variable weight vector. A higher value of $\lambda_{2}$ results in increased sparsity of the variable vector $W$. Specifically, a higher weight value $w_{j}$ indicates greater importance of the $j$th variable in the clustering process. Conversely, when $w_{j}$ is equal to zero, it signifies that the $j$th variable does not contribute to the clustering.
In order to gain a deeper comprehension of the algorithmic process we have proposed, we summarize the aforementioned procedures in a pseudocode in Algorithm 1.
@@
1:Initialize the weight vector $W$ as $w_{j}=1/\sqrt{p}$, set tolerance $\varepsilon>0$ and $r=0$.
@@
5:     Calculate the new $E_{ij}$ by (7) or (8), which incorporates the cluster-specific means $\mu_{k\cdot}=|n_{k}|^{-1}\sum_{j}\sum_{i\in c_{k}}x_{ij}^{*}$
@@
8:     Arrange the cluster $\Theta^{(r+1)}$ and compute the $Q_{j}^{R}(x_{\cdot j};\Theta^{(r+1)},E)$ for different variables by:
$Q_{j}^{R}(x_{\cdot j};\Theta^{(r+1)},E)=\sum_{i=1}^{n}\Big{(}x_{ij}^{\prime}-\frac{1}{n}\sum_{i=1}^{n}x_{ij}^{\prime}\Big{)}^{2}-\sum_{k=1}^{K}\sum_{i\in c^{(r+1)}_{k}}\Big{(}x_{ij}^{\prime}-\frac{1}{|n_{k}|}\sum_{i\in c^{(r+1)}_{k}}x_{ij}^{\prime}\Big{)}^{2},$
@@
$w^{(r+1)}_{j}=\frac{S(Q_{j}^{R}(x_{\cdot j};\Theta,E_{\cdot j});\lambda_{2})}{\sqrt{\sum_{j=1}^{p}(S(Q_{j}^{R}(x_{\cdot j};\Theta,E_{\cdot j});\lambda_{2}))^{2}}}$
@@
There are two tuning parameters in the proposed algorithm: $\lambda_{1}$ and $\lambda_{2}$. The tuning parameter $\lambda_{1}$ is crucial for finding the error matrix and impacts outliers’ detection, and the tuning parameter $\lambda_{2}$ plays a crucial role in filtering out variables contributing to the clustering process. We here provide the robust $\rm Gap$ statistics to determine these parameters.
@@
$\begin{split}{\rm Gap}_{\lambda_{2},\lambda_{1}}=\log(D^{R}_{\lambda_{2},\lambda_{1}})-\frac{1}{B}\sum_{b=1}^{B}\log(^{(b)}D^{R}_{\lambda_{2},\lambda_{1}}),\ \ \ \ D^{R}_{\lambda_{2},\lambda_{1}}=\sum^{p}_{j=1}w_{j}Q_{j}^{R}(x_{\cdot j};\Theta,E_{\cdot j}),\end{split}$ (10)
and the optimal value of $\lambda_{2}$ and $\lambda_{1}$ corresponding to the largest value of ${\rm Gap}_{\lambda_{2},\lambda_{1}}$. Note that $D^{R}_{\lambda_{2},\lambda_{1}}$ is a weighted robust between-cluster sum of squares defined in (3), and ${}^{(b)}D^{R}_{\lambda_{2},\lambda_{1}}$ is a version for permuted datasets. Here $B$ denotes the number of permuted datasets, which are generated by randomly selecting from the original dataset. Increasing the number of permuted datasets leads to improved accuracy in selecting the true parameters. The between-cluster sum of squares, taking account of the variable weight, is adopted in Witten and Tibshirani (2010), so our version used in (10) can be regarded as its robust extension.
Since it may be computationally intensive to search the optimal value of $(\lambda_{2},\lambda_{1})$ by a grid search method, we conduct an alternating optimization algorithm for tuning parameter search. Specifically, we first set a suitable value $\lambda_{1}^{\dagger}$ for $\lambda_{1}$ and compute the optimal value $\lambda_{2}^{\ast}$ of $\lambda_{2}$ by maximizing $\rm Gap_{\lambda_{2},\lambda_{1}^{\dagger}}$. Then, we obtain the optimal value $\lambda_{1}^{\ast}$ of $\lambda_{1}$ by maximizing $\rm Gap_{\lambda_{2}^{\ast},\lambda_{1}}$. Using this search algorithm instead of the grid search method can save a significant huge of computation time. We also offer a pseudocode for this procedure for easy understanding of this search algorithm given in Algorithm 2.
@@
$\begin{split}&{\rm Gap}_{\lambda_{2}}=\log(D^{R}_{\lambda_{2},\lambda_{1}^{\dagger}})-\frac{1}{B}\sum_{b=1}^{B}\log(^{(b)}D^{R}_{\lambda_{2},\lambda_{1}^{\dagger}})\end{split}$
@@
$\begin{split}&{\rm Gap}_{\lambda_{1}}=\log(D^{R}_{\lambda_{2}^{*},\lambda_{1}})-\frac{1}{B}\sum_{b=1}^{B}\log(^{(b)}D^{R}_{\lambda_{2}^{*},\lambda_{1}})\end{split}$
@@
3:Output the optimal set of tuning parameters, $(\lambda_{2}^{\ast},\lambda_{1}^{\ast})$
@@
In this section, we explore the ability of the proposed clustering method. We consider that each observation $x_{i}$ is generated independently from a multivariate normal distribution, given that the observation belongs to cluster $k$. Specifically, for an observation $x_{i}$ in cluster $k$, we have $x_{i}\sim\mathcal{N}(\mu_{k\cdot},\Sigma_{p})$, where $\mu_{k\cdot}\in\mathbb{R}^{p}$ denotes the mean vector for cluster $k$, and $\Sigma_{p}\in\mathbb{R}^{p\times p}$ represents the covariance matrix. Each element of $\mu_{k\cdot}$ is independently sampled from from either $U(-6,-3)$ or $U(3,6)$. To simulate scenarios involving outliers, we introduce a contaminated error distribution with a multivariate normal mixture, represented as $(1-\pi)\mathcal{N}(\mu_{k\cdot},\Sigma_{p})+\pi\mathcal{N}(\mu_{k\cdot}+b_{j},\Sigma_{p}),j=1,\ldots,p$ where $\pi\in[0,1]$ represents the proportion of outliers for each cluster. In the simulation study, We consider two types of $\Sigma_{p}$. In the first scenario, the assumption is made that all variables are independent. In another scenario, we assume there is a correlation structure among variables. To introduce randomness to the appearance of outliers, $b_{j}$ is generated with a random number from one of two uniform distributions: $U(-13,-7)$ or $U(7,13)$. In this setup, we assume that some explanatory variables are significantly associated with the response, while the remaining variables are redundant for variable selection. To mimic this situation, we randomly set $q$ elements of $\mu_{k\cdot}$ to non-zero values while the remaining elements are set to zero. Consequently, a successful method should accurately identify the $q$ positions where the weights $w_{j}$ corresponding to the true significant variables are non-zero while ensuring that the weights $w_{j}$ of the remaining variables are set to zero.
To evaluate the approach’s clustering accuracy capability, we employ the clustering error rate (CER) (Rand, 1971). The CER measures the extent to which the model’s predicted partition ${\widehat{C}}$ for a set of $n$ observations is consistent with the true labels $C$. The CER is defined as
${\rm CER}({\widehat{C}},C)=\binom{n}{2}^{-1}\sum_{i<i^{\prime}}\big{|}I_{{\widehat{C}}_{(i,i^{\prime})}}-I_{C_{(i,i^{\prime})}}\big{|},$
where, $I_{C_{(i,i^{\prime})}}$ and $I_{{\widehat{C}}_{(i,i^{\prime})}}$ indicate whether the $i$th and $i^{\prime}$th observations belong to the same cluster. Lower CER indicates better performance in clustering accuracy or outlier detection.
@@
In this subsection explores the new $\rm Gap$ statistics’ capability to directly search for the tuning parameters of robustness and variable vector sparsity using Algorithm 2. Initially, we considered 3 clusters, each containing 50 observations, number of variables as $p=50$, with $q=5$, and all variables are independent, i.e., $\Sigma_{p}=I_{p}$. A proper of hyperparameters $\lambda_{1}$ and $\lambda_{2}$ should enable the model to accurately identify the number of outliers and the number of informative variables in different contamination levels for $\pi$ in $\left\{0,0.1,0.2,0.3\right\}$. The proposed tuning parameter search strategy generates both the robustness parameter $\lambda_{1}$ and the variable selection tuning parameter $\lambda_{2}$ using an exponential decay. We consider the number of permuted datasets to be 25 (i.e., $B=25$).We employed different thresholding function types, where the first thresholding function is used to control model sparsity, and the second thresholding function is used to control model robustness. The tuning parameter search process is repeated 30 times for 30 different datasets, and the resulting table is presented in Table 1.
@@
Table 1 summarizes the results of the evaluation measures. Overall, a good tuning parameter $\lambda_{1}$ should accurately identify the number of outliers under different contamination levels. As the contamination level $\pi$ increases from 0 to 0.3, we observe that the four thresholding methods demonstrate varying outlier detection capabilities across different contamination levels. At $\pi=0.1$, the detected number of outliers ranges from 13.33 to 15.56, with soft-soft-thresholding achieving the highest average of 15.56 (standard error: 6.425). When $\pi$ increases to 0.2, the detected number of outliers falls between 24.10 and 28.73, with soft-soft and soft-SCAD-thresholding showing similar performance. This demonstrates that the Gap statistics can help the ARKC method determine $\lambda_{1}$ to some extent. However, when the contamination level reaches 0.3, the soft-soft-thresholding ARSK method detects 33.10 outliers on average, which is lower than the SCAD-SCAD-thresholding ARSK method (44.00 outliers) and deviates from the true number of outliers. We consider that the $L_{1}$ norm penalty has lower robustness compared to the SCAD penalty (Hampel et al. (1986)). Regarding the selection of the tuning parameter $\lambda_{2}$, a good tuning parameter should accurately identify the number of informative variables under different contamination levels. When the data is clean ($\pi=0$), all four methods identify around 4 to 5 informative variables, which is close to 5. As the contamination level increases, the number of detected informative variables generally increases, with some fluctuations. At $\pi=0.3$, the soft-soft-thresholding ARSK method detects 13.60 informative variables on average, which significantly deviates from the 5. In contrast, the SCAD-SCAD-thresholding ARSK method detects 7.187 informative variables, showing better performance in the presence of high contamination. Considering these results, we can conclude that the proposed parameter selection algorithm based on the Gap statistic works relatively well under low to moderate contamination levels. However, its performance may degrade when the contamination level is high, especially for the soft-thresholding-based methods. The SCAD-thresholding-based methods generally show better robustness and accuracy in identifying outliers and informative variables under various contamination levels.
Simulation 2: comparison with other approaches
In this subsection, we present the findings of a comprehensive simulation study conducted to investigate the cluster data structure and properties by the ARSK algorithm based on soft-thresholding and SCAD-thresholding. The study extensively applied $\rm Gap$ statistics to estimate optimal parameter settings and compared with several benchmark approaches, including $K$-means (KC), PCA-$K$-means (PCA-KC) clustering using only the first principal component, trimmed $K$-means (TKM), robust and sparse $K$-means RSKC ($\alpha=0.1,0.2$) (Kondo et al., 2016), and the weighted robust $K$-means (WRCK) algorithms (Brodinova et al., 2019). Note that WRCK and RSKC are both trimming-based clustering approaches.
@@
Considering the mixture error model mentioned at the beginning, we also generated data for 3 clusters, each containing 50 observations. We set the number of variables as $p=50$, with $q=5$, and $p=500$, with $q=50$. Meanwhile, we consider $\pi$ in $\left\{0,0.1,0.2\right\}$ for specifying the distribution of outliers. As previously stated, we study two types of covariance matrices: one with $\Sigma_{p}=I_{p}$ and the another generated according to the method proposed by Hirose et al. (2017) as
$\Sigma_{p}=Q\begin{bmatrix}1&\rho_{t}&\cdots&\rho_{t}\\ \rho_{t}&1&\ddots&\rho_{t}\\ \vdots&\ddots&1&\vdots\\ \rho_{t}&\cdots&\rho_{t}&1\end{bmatrix}Q^{T},$
@@
where the $Q$ denotes a $p\times p$ random rotation matrix satisfying $Q^{T}=Q^{-1}$ and the $\rho_{t}$ are randomly generated form an Uniform distribution $U(0.1,1)$.
@@
In order to evaluate the clustering accuracy and outlier detection capability of each approach, we employ the clustering error rate (CER) as described above. To assess the outlier detection performance, the outliers identified by each model are assigned to the $(K+1)$-th cluster group. Table 2 presents the results of 100 simulations for the scenario where all variables are independent, i.e., $\Sigma_{p}=I_{p}$, while Table 3 summarizes the results of 100 simulations for the scenario where variables exhibit a certain correlation structure.
@@
In Tables 2 and 3, the results show that KC, PCA-KC, and TKM approaches, which lack robustness for high dimensions or outliers, are the worst performers, particularly when the dimensionality $p$ and proportion of outliers $\pi$ increase. For instance, in the independent case with $p=500$ and $q=50$, the CER of KC increases from 0.050 to 0.341 as $\pi$ increases from 0 to 0.2. Similarly, the CER of PCA-KC and TKM ($\alpha=0.1$) increase from 0.058 to 0.300 and 0.128 to 0.302, respectively, under the same scenario. Conversely, RSCK, WRCK, and ARSK perform exceptionally well in high-dimension and high-contamination scenarios. However, the effectiveness of the RSCK approach heavily depends on the choice of the parameter $\alpha$. When $\alpha$ equals the true outlier proportion, RSCK performs impressively, with CER close to 0. For example, in the correlated case with $p=500$ and $q=50$, the CER of RSKC ($\alpha=0.1$) is 0 when $\pi=0.1$. However, the approach loses effectiveness if $\alpha$ is incorrectly specified, leading to higher CER values. In the same scenario, the CER of RSKC ($\alpha=0.2$) is 0.122 when $\pi=0.1$. The WRCK approach, based on the density clustering method, generally outperforms RSCK when outliers are present. Still, it may erroneously classify more normal data points as outliers when no contamination exists. In contrast, the ARSK approach, with different combinations of thresholding functions (e.g., soft-soft, soft-SCAD, SCAD-soft and SCAD-SCAD), consistently maintains a low Clustering Error Rate (CER) across all simulated data scenarios.
In summary, the simulation results demonstrate the efficacy and robustness of the Adaptive Robust Sparse K-means (ARSK) method in clustering high-dimensional datasets, particularly in the presence of outliers and variable correlations. Across diverse data scenarios, the ARSK variants consistently outperform traditional methods, including K-means (KC), PCA-K-means (PCA-KC), Trimmed K-means (TKM), Robust and Sparse K-means (RSKC), and Weighted Robust K-means (WRCK), highlighting the distinct advantages of the ARSK approach.
@@
Through 100 repetitions, the results of the TPR and TNR are presented in Tables 4 and 5. When variables are independent (Table 4), the TPR and TNR of all methods are generally higher compared to the correlated case (Table 5). This indicates that when variables are correlated with each other, all models tend to make less accurate judgments about informative and non-informative variables, emphasizing the challenges posed by variable correlations in the context of clustering and variable selection.
As the proportion of outliers increases from 0 to 0.2, both the TPR and TNR of most approaches exhibit a decreasing trend across different scenarios, suggesting that the presence of outliers can negatively influence the variable selection performance. However, the TPR and TNR of each type of ARSK approach variant remain relatively high and stable across various data scenarios, demonstrating ARSK’s superior variable selection capabilities in the presence of outlier contamination in datasets. For example, in the independent case with $p=500$ and $q=50$, the TPR of soft-soft-ARSK, soft-SCAD-ARSK, SCAD-soft-ARSK, and SCAD-SCAD-ARSK remain above 0.797, 0.810, 0.798, and 0.810, respectively, even when $\pi=0.2$. Similarly, their TNR values stay above 0.977, 0.980, 0.978, and 0.981 in the same scenario. These results emphasize the superior ability of soft-ARSK and SCAD-ARSK methods to accurately identify informative and non-informative variables in the presence of outliers, making them valuable tools for handling contaminated clustering datasets.

### v2

@@ Introduction
When applying the $P_{2}(w_{j};\lambda_{2})=\lambda_{2}|w_{j}|$ penalty in (9), $S$ can be represented by the soft-thresholding operator as
@@
When applying the $P_{2}(w_{j};\lambda_{2})=P^{SCAD}_{\lambda_{2}}(w_{j})$ penalty in (9), $S$ can be represented by the SCAD-thresholding operator as
@@
The parameter $\lambda_{2}$ has a crucial role in controlling the sparsity in the variable weight vector. A higher value of $\lambda_{2}$ results in increased sparsity of the variable vector $\bm{w}$. Specifically, a higher weight value $w_{j}$ indicates greater importance of the $j$th variable in the clustering process. Conversely, when $w_{j}$ is equal to zero, it signifies that the $j$th variable does not contribute to the clustering.
In order to gain a deeper comprehension of the algorithmic process we have proposed, we summarize the aforementioned procedures in a pseudocode in Algorithm 1. Moreover, the computation complexity of $T$ iteration of Algorithm 1 can be represented $O(nptTK)$, assuming that the number of iterations in the 4th step is $O(t)$.
@@
1:Initialize the weight vector $\bm{w}$ as $w_{j}=1/\sqrt{p}$, set tolerance $\varepsilon>0$ and $r=0$.
@@
5:     Calculate the new $E_{ij}$ by (7) or (8), which incorporates the cluster-specific means $\bm{\mu}_{k,:}=|n_{k}|^{-1}\sum_{j}\sum_{i\in c_{k}}X_{ij}^{*}$
@@
8:     Arrange the cluster $\Theta^{(r+1)}$ and compute the $Q_{j}^{R}(\bm{X}_{:,j};\Theta^{(r+1)},\bm{E})$ for different variables by:
$Q_{j}^{R}(\bm{X}_{:,j};\Theta^{(r+1)},\bm{E})=\sum_{i=1}^{n}\Big{(}X_{ij}^{\prime}-\frac{1}{n}\sum_{i=1}^{n}X_{ij}^{\prime}\Big{)}^{2}-\sum_{k=1}^{K}\sum_{i\in c^{(r+1)}_{k}}\Big{(}X_{ij}^{\prime}-\frac{1}{|n_{k}|}\sum_{i\in c^{(r+1)}_{k}}X_{ij}^{\prime}\Big{)}^{2},$
@@
$w^{(r+1)}_{j}=\frac{S(Q_{j}^{R}(\bm{X}_{:,j};\Theta,\bm{E}_{:,j});\lambda_{2})}{\sqrt{\sum_{j=1}^{p}(S(Q_{j}^{R}(\bm{X}_{:,j};\Theta,\bm{E}_{:,j});\lambda_{2}))^{2}}}$
@@
There are two tuning parameters in the proposed algorithm: $\lambda_{1}$ and $\lambda_{2}$. The tuning parameter $\lambda_{1}$ is crucial for finding the error matrix and impacts the detection of outliers, and the tuning parameter $\lambda_{2}$ plays a crucial role in filtering out variables contributing to the clustering process. We here provide the robust $\rm Gap$ statistics to determine these parameters.
@@
$\begin{split}{\rm Gap}_{\lambda_{2},\lambda_{1}}=\log(D^{R}_{\lambda_{2},\lambda_{1}})-\frac{1}{B}\sum_{b=1}^{B}\log(D^{R(b)}_{\lambda_{2},\lambda_{1}}),\ \ \ \ D^{R}_{\lambda_{2},\lambda_{1}}=\sum^{p}_{j=1}w_{j}Q_{j}^{R}(\bm{X}_{:,j};\Theta,\bm{E}_{:,j}),\end{split}$ (10)
and the optimal value of $\lambda_{2}$ and $\lambda_{1}$ corresponding to the largest value of ${\rm Gap}_{\lambda_{2},\lambda_{1}}$. Note that $D^{R}_{\lambda_{2},\lambda_{1}}$ is a weighted robust between-cluster sum of squares defined in (3), and $D^{R(b)}_{\lambda_{2},\lambda_{1}}$ is a version with $b$th permuted datasets. Here, $B$ denotes the number of permuted datasets generated by randomly selecting from the original dataset. Increasing the number of permuted datasets improves the accuracy of selecting the true parameters. The between-cluster sum of squares, taking account of the variable weight, is adopted in Witten and Tibshirani (2010), so our version used in (10) can be regarded as its robust extension.
Since it may be computationally intensive to search the optimal value of $(\lambda_{2},\lambda_{1})$ by a grid search method, we conduct an alternating optimization algorithm for tuning parameter search. Specifically, we first set a suitable value $\lambda_{1}^{\dagger}$ for $\lambda_{1}$ and compute the optimal value $\lambda_{2}^{\ast}$ of $\lambda_{2}$ by maximizing $\rm Gap_{\lambda_{2},\lambda_{1}^{\dagger}}$. Then, we obtain the optimal value $\lambda_{1}^{\ast}$ of $\lambda_{1}$ by maximizing $\rm Gap_{\lambda_{2}^{\ast},\lambda_{1}}$. Using this search algorithm instead of the grid search method can save a significant computation time. We also offer a pseudocode for this procedure for easy understanding of this search algorithm given in Algorithm 2.
@@
$\begin{split}&{\rm Gap}_{\lambda_{2}}=\log(D^{R}_{\lambda_{2},\lambda_{1}^{\dagger}})-\frac{1}{B}\sum_{b=1}^{B}\log(D^{R(b)}_{\lambda_{2},\lambda_{1}^{\dagger}})\end{split}$
@@
$\begin{split}&{\rm Gap}_{\lambda_{1}}=\log(D^{R}_{\lambda_{2}^{*},\lambda_{1}})-\frac{1}{B}\sum_{b=1}^{B}\log(D^{R(b)}_{\lambda_{2}^{*},\lambda_{1}})\end{split}$
@@
3:Output: the optimal set of tuning parameters, $(\lambda_{2}^{\ast},\lambda_{1}^{\ast})$
@@
Experimental setup
In this section, we explore the ability of the proposed clustering method. We consider that each observation $\bm{x}_{i}$ is generated independently from a multivariate normal distribution, given that the observation belongs to cluster $k$. Specifically, for an observation $\bm{x}_{i}$ in cluster $k$, we have $\bm{x}_{i}\sim\mathcal{N}(\bm{\mu}_{k,:},\bm{\Sigma}_{p})$, where $\bm{\mu}_{k,:}\in\mathbb{R}^{p}$ denotes the mean vector for cluster $k$, and $\bm{\Sigma}_{p}\in\mathbb{R}^{p\times p}$ represents the covariance matrix. Each element of $\bm{\mu}_{k,:}$ is independently sampled from from either $U(-6,-3)$ or $U(3,6)$. To simulate scenarios involving outliers, we introduce a contaminated error distribution with a multivariate normal mixture, represented as $(1-\pi)\mathcal{N}(\bm{\mu}_{k,:},\bm{\Sigma}_{p})+\pi\mathcal{N}(\bm{\mu}_{k,:}+b_{j},\bm{\Sigma}_{p})$ where $j\in(1,\dots,p)$ and $\pi\in[0,1]$ represents the proportion of outliers for each cluster. In the simulation study, we consider two types of $\bm{\Sigma}_{p}$. In the first scenario, the assumption is made that all variables are independent. In another scenario, we assume there is a correlation structure among variables. To introduce randomness to the appearance of outliers, $b_{j}$ is generated with a random number from one of two uniform distributions: $U(-13,-7)$ or $U(7,13)$. In this setup, we assume that some explanatory variables are significantly associated with the response, while the remaining variables are redundant for variable selection. To mimic this situation, we randomly set $q$ elements of $\bm{\mu}_{k,:}$ to non-zero values while the remaining elements are set to zero. Consequently, a successful method should accurately identify the $q$ positions where the weights $w_{j}$ corresponding to the true significant variables are non-zero while ensuring that the weights $w_{j}$ of the remaining variables are set to zero.
In order to evaluate the approach’s clustering accuracy capability, we employ the clustering error rate (CER) (Rand, 1971). The CER measures the extent to which the model’s predicted partition $\hat{C}$ for a set of $n$ observations is consistent with the true labels $C$.
The CER is defined as
${\rm CER}(\hat{C},C)=\binom{n}{2}^{-1}\sum_{i<i^{\prime}}\big{|}1_{\hat{C}_{(i,i^{\prime})}}-1_{C_{(i,i^{\prime})}}\big{|},$ (11)
where, $1_{C_{(i,i^{\prime})}}$ and $1_{\hat{C}_{(i,i^{\prime})}}$ indicate whether the $i$th and $i^{\prime}$th observations belong to the same cluster. Lower CER indicates better performance in clustering accuracy or outlier detection.
@@
In this subsection explores the new $\rm Gap$ statistics’ capability to directly search for the tuning parameters of robustness and variable vector sparsity using Algorithm 2. Initially, we considered 3 clusters, each containing 50 observations, number of variables as $p=50$, with $q=5$, and all variables are independent, i.e., $\bm{\Sigma}_{p}=\bm{I}_{p}$. A proper of hyperparameters $\lambda_{1}$ and $\lambda_{2}$ should make the model accurately identify the structure of the clustering data in different contamination levels for $\pi$ in $\left\{0,0.1,0.2,0.3\right\}$. The proposed tuning parameter search strategy generates both the robustness parameter $\lambda_{1}$ and the variable selection tuning parameter $\lambda_{2}$ using exponential decay. We consider the number of permuted datasets to be 25 (i.e., $B=25$). We employed different thresholding function types, where the first thresholding function is used to control model sparsity, and the second thresholding function is used to control model robustness. The tuning parameter search process is repeated 30 times for 30 different datasets, and the resulting table is presented in Table 1.
@@
Table 1 summarizes the results of the evaluation measures. Overall, a good tuning parameter $\lambda_{1}$ should accurately identify the number of outliers under different contamination levels. As the contamination level $\pi$ increases from 0 to 0.3, we observe that the four thresholding methods demonstrate varying outlier detection capabilities across different contamination levels. At $\pi=0.1$, the detected number of outliers ranges from 13.33 to 15.56, with soft-soft-thresholding achieving the highest average of 15.56 (standard error: 6.425). When $\pi$ increases to 0.2, the detected number of outliers falls between 24.10 and 28.73, with soft-soft and soft-SCAD-thresholding showing similar performance. This demonstrates that the Gap statistics can help the ARKC method determine $\lambda_{1}$ to some extent. However, when the contamination level reaches 0.3, the soft-soft-thresholding ARSK method detects 33.10 outliers on average, which is lower than the SCAD-SCAD-thresholding ARSK method (44.00 outliers) and deviates from the true number of outliers.
Regarding the selection of the tuning parameter $\lambda_{2}$, a good tuning parameter should accurately identify the number of informative variables under different contamination levels. When the data is clean ($\pi=0$), all four methods identify around 4 to 5 informative variables, which is close to 5. As the contamination level increases, the number of detected informative variables generally increases, with some fluctuations. At $\pi=0.3$, the soft-soft-thresholding ARSK method detects 13.60 informative variables on average, which significantly deviates from the 5. In contrast, the SCAD-SCAD-thresholding ARSK method detects 7.187 informative variables, showing better performance in the presence of high contamination. Furthermore, errors in identifying informative variables can lead to significant biases in outlier detection.
Considering these results, we can conclude that the proposed parameter selection algorithm based on the Gap statistic works relatively well under low to moderate contamination levels. However, its performance may degrade when the contamination level is high, especially for the soft-thresholding-based methods. The SCAD-thresholding-based methods generally show better robustness and accuracy in identifying outliers and informative variables under various contamination levels.
Simulation 2: comparison with other methods
In this subsection, we present the findings of a comprehensive simulation study conducted to investigate the cluster data structure and properties by the ARSK algorithm based on soft-thresholding and SCAD-thresholding. The study extensively applied $\rm Gap$ statistics to estimate optimal parameter settings and compared with several benchmark approaches, including the original $K$-means (KC), PCA-$K$-means (PCA-KC) - a mixed approach that combines Principal Component Analysis (PCA) with the original $K$-means algorithm, as initially proposed by Ding and He (2004). It is noted that PCA is a widely recognized dimensionality reduction technique that transforms high-dimensional data into a lower-dimensional space. We also considered Trimmed $K$-means (TKM), which executes standard $K$-means clustering and subsequently eliminates a predetermined percentage of points exhibiting the greatest Euclidean distance from their respective cluster centroids. Furthermore, we employ Robust and Sparse $K$-means (RSKC), whose fundamental premise involves adapting the SK-means algorithm of Witten and Tibshirani (2010) in conjunction with Trimmed $K$-means. Lastly, we apply the Weighted Robust $K$-means (WRCK) method, which shares similarities with RSKC but distinguishes itself through the incorporation of observation-specific weights in its objective function. These weights are reflective of each observation’s isolation level relative to its surrounding neighbourhood.
@@
Considering the mixture error model mentioned at the beginning, we also generated data for 3 clusters, each containing 50 observations. We set the number of variables as $p=50$, with $q=5$, and $p=500$, with $q=50$. Meanwhile, we consider $\pi$ in $\left\{0,0.1,0.2\right\}$ for specifying the distribution of outliers. As previously stated, we study two types of covariance matrices: one with $\bm{\Sigma}_{p}=\bm{I}_{p}$ and the another generated according to the method proposed by Hirose et al. (2017) as
$\bm{\Sigma}_{p}=\bm{Q}\begin{bmatrix}1&\rho_{t}&\cdots&\rho_{t}\\ \rho_{t}&1&\ddots&\rho_{t}\\ \vdots&\ddots&1&\vdots\\ \rho_{t}&\cdots&\rho_{t}&1\end{bmatrix}\bm{Q}^{T},$
@@
where the $\bm{Q}$ denotes a $p\times p$ random rotation matrix satisfying $\bm{Q}^{T}=\bm{Q}^{-1}$ and the $\rho_{t}$ are randomly generated form an Uniform distribution $U(0.1,1)$.
@@
In order to evaluate the clustering accuracy and outlier detection capability of each approach, we employ the clustering error rate (CER) as described above. To assess the outlier detection performance, the outliers identified by each model are assigned to the $(K+1)$-th cluster group. Table 2 presents the results of 100 simulations for the scenario where all variables are independent, i.e., $\bm{\Sigma}_{p}=\bm{I}_{p}$, while Table 3 summarizes the results of 100 simulations for the scenario where variables exhibit a certain correlation structure.
@@
In Tables 2 and 3, the results show that KC, PCA-KC, and TKM approaches, which lack robustness for high dimensions or outliers, are the worst performers, particularly when the dimensionality $p$ and proportion of outliers $\pi$ increase. For instance, in the independent case with $p=500$ and $q=50$, the CER of KC increases from 0.050 to 0.341 as $\pi$ increases from 0 to 0.2. Similarly, the CER of PCA-KC and TKM ($\alpha=0.1$) increase from 0.058 to 0.300 and 0.128 to 0.302, respectively, under the same scenario. Conversely, RSCK, WRCK, and ARSK perform exceptionally well in high-dimension and high-contamination scenarios. However, the effectiveness of the RSCK approach heavily depends on the choice of the parameter $\alpha$. When $\alpha$ equals the true outlier proportion, RSCK performs impressively, with CER close to 0. For example, in the correlated case with $p=500$ and $q=50$, the CER of RSKC ($\alpha=0.1$) is 0 when $\pi=0.1$. However, the approach loses effectiveness if $\alpha$ is incorrectly specified, leading to higher CER values. In the same scenario, the CER of RSKC ($\alpha=0.2$) is 0.122 when $\pi=0.1$. The WRCK approach, based on the density clustering method, generally outperforms RSCK when outliers are present. Still, it may erroneously classify more normal data points as outliers when no contamination exists. In contrast, the ARSK approach, with different combinations of thresholding functions (e.g., soft-soft, soft-SCAD, SCAD-soft, and SCAD-SCAD), consistently maintains a low Clustering Error Rate (CER) across all simulated data scenarios.
In summary, the simulation results demonstrate the effectiveness and robustness of the ARSK method in clustering high-dimensional datasets, particularly in the presence of outliers and variable correlations. Across diverse data scenarios, the ARSK variants consistently outperform traditional methods, including KC, PCA-KC, TKM, RSKC, and WRCK, highlighting the distinct advantages of the ARSK approach.
@@
Through 100 repetitions, the results of the TPR and TNR are presented in Tables 4 and 5. As the proportion of outliers increases from 0 to 0.2, both the TPR and TNR of most approaches demonstrate a decreasing trend across different scenarios, suggesting that the presence of outliers can negatively influence the variable selection performance. However, the TPR and TNR of each type of ARSK approach variant remain relatively high and stable across various data scenarios, demonstrating ARSK’s superior variable selection capabilities in the presence of outlier contamination in datasets. For example, in the independent case with $p=500$ and $q=50$, the TPR of soft-soft-ARSK, soft-SCAD-ARSK, SCAD-soft-ARSK, and SCAD-SCAD-ARSK remain above 0.797, 0.810, 0.798, and 0.810, respectively, even when $\pi=0.2$. Similarly, their TNR values stay above 0.977, 0.980, 0.978, and 0.981 in the same scenario. These results emphasize the superior ability of soft-ARSK and SCAD-ARSK methods to accurately identify informative and non-informative variables in the presence of outliers, making them valuable tools for handling contaminated clustering datasets.